OpenStudy (anonymous):

Find the 15th term in the geometric sequence below. 20, 10, 5, 2.5, . . .

OpenStudy (gorv):

common ratio= 10/20=1/2

OpenStudy (gorv):

t20= a*r^(20-1)=a*r^19

OpenStudy (gorv):

a=20 r=1/2